# Statistical Test on Diffusion Model-based Anomaly Detection by Selective Inference

### v3

@@ Abstract
Advancements in AI image generation, particularly diffusion models, have progressed rapidly. However, the absence of an established framework for quantifying the reliability of AI-generated images hinders their use in critical decision-making tasks, such as medical image diagnosis. In this study, we address the task of detecting anomalous regions in medical images using diffusion models and propose a statistical method to quantify the reliability of the detected anomalies. The core concept of our method involves a selective inference framework, wherein statistical tests are conducted under the condition that the images are produced by a diffusion model. With our approach, the statistical significance of anomaly detection results can be quantified in the form of a $p$-value, enabling decision-making with controlled error rates, as is standard in medical practice. We demonstrate the theoretical soundness and practical effectiveness of our statistical test through numerical experiments on both synthetic and brain image datasets.
@@ 1 Introduction
Advances in image generation AI, such as diffusion models, have been remarkable (Song and Ermon, 2019). They can generate highly realistic and detailed images, which leads to innovations in various tasks across various fields. For example, image generation AI significantly enhances medical image diagnosis by improving accuracy and efficiency. It can generate highly detailed and enhanced images from standard medical scan images, potentially offering doctors to detect anomalies and diseases with greater precision. Furthermore, image generation AI can be used to create alternative versions of medical images to consider what-if scenarios. For example, it can generate virtual images of a patient when they are healthy, which allows for comparing the current actual images with the virtual healthy images, making it possible to provide a diagnosis tailored to the individual patient.
On the other hand, when using virtual images generated by AI for critical decision-making, such as medical diagnosis, it is crucial to ensure the reliability of the decisions. Given that images are generated by an AI algorithm, such as a deep learning model trained on historical data, they may inherently contain biases and errors. Therefore, treating virtual synthetic images as equivalent to real images in decision-making tasks carries the risk of biased and erroneous outcomes. When making critical decisions based on generated images, it is necessary to be able to assess their reliability by properly taking into account the fact that the images were generated by AI. However, to our knowledge, there are no studies that can quantify the reliability of decision-making based on image generation AI.
Figure 1: Schematic illustration of the anomaly detection on a brain image dataset using a diffusion model and the proposed DMAD-test. When a test image, which may contain an anomalous region, is fed into a trained diffusion model, a normal image is generated through the forward process and reverse process. By initiating the image generation from the middle of the forward process, a normal image that retains the characteristics of the input image can be generated. By comparing the input image with the normal image, the anomalous region can be identified. In this study, we propose a method called the DMAD-test, which quantifies the statistical significance of the identified anomalous regions in the form of $p$-value. The DMAD-test calculates the $p$-values by incorporating the fact that the anomalous region has been identified by the diffusion model, thus enabling unbiased decision-making (see §3 and §4 for details).
In this study, we address this challenge using the statistical hypothesis testing framework. We introduce a statistical inference framework called selective inference (SI), which has gained attention over the past decade in the statistics community as a novel approach for data-driven hypotheses (Taylor and Tibshirani, 2015; Fithian et al., 2015; Lee and Taylor, 2014). In SI, statistical inference is performed based on the sampling distribution of the test statistic under the condition that the hypothesis being tested was selected based on the data. Our core idea is to formulate decision-making tasks involving generated images as statistical hypothesis testing problems, and to incorporate SI framework to accurately quantify the reliability of decisions influenced by these generated images.
As an example of decision-making tasks based on image generation AI, we focus on the problem of detecting anomalous regions in medical images (Wolleb et al., 2022; Baur et al., 2021) (see Figure 1). Initially, a diffusion model is trained exclusively on normal images during the training phase. In the testing phase, a patient’s test image is processed through this model to create a virtual normal image, against which the original is compared to identify anomalous regions. Our proposed statistical test, the Diffusion Model-based Anomalous Region Detection Test (DMAD-test), quantifies the statistical reliability of detected anomalies as $p$-values. Decisions based on these $p$-values can theoretically control the false detection rate at desired significance levels (such as 0.01 or 0.05).
Related work.
@@
Diffusion models have been effectively utilized in anomalous region detection problems (Wolleb et al., 2022; Pinaya et al., 2022; Fontanella et al., 2023; Wyatt et al., 2022; Mousakhan et al., 2023). In this context, the denoising diffusion probabilistic model (DDPM) is commonly used (Ho et al., 2020; Song et al., 2022). During the training phase, a DDPM model learns the distribution of normal medical images by iteratively adding and then removing noise. In the test phase, the model attempts to reconstruct a new test image. If the image contains anomalous regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. The discrepancies between the original and the reconstructed image are then analyzed to identify and highlight anomalous regions. Other types of generative AI has also been used for anomalous region detection task (Baur et al., 2021; Chen and Konukoglu, 2018; Chow et al., 2020; Jana et al., 2022).
@@
SI was first introduced within the context of reliability evaluation for linear model features when they were selected using a feature selection algorithm (Lee and Taylor, 2014; Lee et al., 2016; Tibshirani et al., 2016), and then extended to more complex feature selection methods (Yang et al., 2016; Suzumura et al., 2017; Hyun et al., 2018; Rügamer and Greven, 2020; Das et al., 2021). Then, SI proves valuable not only for feature selection problems but also for statistical inference across various data-driven hypotheses, including unsupervised learning tasks (Chen and Bien, 2020; Tsukurimichi et al., 2021; Tanizaki et al., 2020; Duy et al., 2022; Le Duy et al., 2024; Lee et al., 2015; Gao et al., 2022; Duy et al., 2020; Jewell et al., 2022). The fundamental idea of SI is to perform an inference conditional on the hypothesis selection event, which mitigates the selection bias issue even when the hypothesis is selected and tested using the same data. To conduct SI, it is necessary to derive the sampling distribution of test statistic conditional on the hypothesis selection event. To the best of our knowledge, SI was applied to statistical inferences on several deep learning models (Duy et al., 2022; Miwa et al., 2023; Shiraishi et al., 2024; Miwa et al., 2024), but none of them works on image generation by diffusion models.
Contributions.
Our main contributions in this study are summarized as follows. Our first contribution is the introduction of a statistical testing framework for quantifying reliability in decision-making based on images generated by diffusion models. The second contribution is the implementation of SI for diffusion models, which requires the calculation of the sampling distribution conditional on the diffusion model, necessitating the development of non-trivial computational methodology. The third contribution is to theoretically guarantee the performance of the proposed DMAD-test and demonstrate its performance through numerical experiments and applications in brain imaging diagnostics. The code is available as supplementary material.
2 Diffusion Models
In this section, we briefly explain the diffusion model employed in this study. Given a test image which possibly contain anomalous regions, a denoising diffusion model (Ho et al., 2020; Song et al., 2022) is used to generate the corresponding normal image. The reconstruction process consists of two processes called forward process (or diffusion process) and reverse process.
@@ 2 Diffusion Models
In the forward process, noise is sequentially added to the test image so that it converges to a standard Gaussian distribution $\mathcal{N}(\bm{0},I)$. Let $\mathbf{x}$ be an image represented as a vector with each element corresponding to a pixel value. Given an original test image $\mathbf{x}_{0}$, noisy images $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ are sequentially generated, where $T$ is the number of noise addition steps. We consider the distribution of the original and noisy test images, which is denoted by $q(\mathbf{x})$, and approximate the distribution by a parametric model $p_{\theta}(\mathbf{x})$ with $\theta$ being the parameters. Using a sequence of noise scheduling parameters $0<\beta_{1}<\beta_{2},<\dots<\beta_{T}<1$, the forward process is written as
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0}):=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),\quad\text{where}\;\;q(\mathbf{x}_{t}|\mathbf{x}_{t-1}):=\mathcal{N}(\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}I).$
By the reproducibility of the Gaussian distribution, $\mathbf{x}_{t}$ can be rewritten by a linear combination of $\mathbf{x}_{0}$ and $\epsilon$, i.e.,
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\epsilon,\;\;\;\text{with}\;\;\;\epsilon\sim\mathcal{N}(\mathbf{0},I),$ (1)
where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$.
@@
In the reverse process, a parametric model in the form of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\beta_{t}I)$ is employed, where $\mu_{\theta}(\mathbf{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$. Typically, a U-Net is used as the model architecture for $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$. In DDPM (Ho et al., 2020), the loss function for training the noise component is simply written as $||\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})-\epsilon_{t}||^{2}_{2}$. Based on (1), given a noisy image $\mathbf{x}_{t}$ after $t$ steps, the reconstruction of the image in the previous step $\mathbf{x}_{t-1}$ is obtained as
$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\mathbf{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})+\sigma_{t}\epsilon_{t},$ (2)
@@
$f^{(t)}_{\theta}(\mathbf{x}_{t}):=(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\mathbf{x}_{t}))/\sqrt{\alpha_{t}},$ (3)
@@
Here, $\eta$ is a hyperparameter that controls the randomness in the reverse process. By setting $\eta=1$, we can create new images by stochastic sampling. On the other hand, if we set $\eta=0$, deterministic sampling is used for image generation. By recursively sampling as in (2), we can obtain a reconstructed image of the original input $\mathbf{x}_{0}$.
In practice, the reverse process starts from $\mathbf{x}_{T^{\prime}}$ with $T^{\prime}<T$. Namely, we reconstruct the original input image not from the completely noisy one, but from a one which still contain individual information of the original input image. The smaller $T^{\prime}$ ensures that the reconstructed image preserves fine details of the input image. Conversely, the larger $T^{\prime}$ results in the retention of only large scale features, thereby converting more of the anomalous regions into normal regions (Ho et al., 2020; Mousakhan et al., 2023). Therefore, $T^{\prime}$ should be set to balance the feature retention of the input image and the conversion of the anomalous region to the normal region. Note that setting $T^{\prime}$ smaller than $T$ has advantages in terms of computational cost. For the purpose of reducing computational cost, various methods have been proposed. For example, one way is to sample while skipping portions of the sampling trajectory (see Appendix A). The image reconstruction scheme by DDPM is summarized in Algorithm 1.
@@
0:  Input image $\mathbf{x}$
1:  $\mathbf{x}_{T^{\prime}}\leftarrow\sqrt{\alpha_{T^{\prime}}}\mathbf{x}+\sqrt{1-\alpha_{T^{\prime}}}\epsilon$
@@
3:     $f^{(t)}_{\theta}(\mathbf{x}_{t})\leftarrow(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\mathbf{x}_{t}))/\sqrt{\alpha_{t}}$
4:     $\mathbf{x}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\mathbf{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})+\sigma_{t}\epsilon_{t}$
@@
5:  Reconstructed image $\mathbf{x}_{0}$
3 Statistical Test on Generated Images by Diffusion Models
In this section, we formulate the statistical test for detecting anomalous regions using images generated by a trained DDPM model. As shown in Figure 1, anomalous region detection by diffusion models is conducted as follows. First, in the training phase, the diffusion model is trained only on normal images. Then, in the test phase, we feed a test image which might contain anomalous regions into the trained diffusion model, and reconstruct it back from a noisy image $\mathbf{x}_{T^{\prime}}$ at step $T^{\prime}<T$. By appropriately selecting $T^{\prime}$, we can generate a normal image that retain individual characteristics of the test input image. If the image does not contain anomalous regions, the reconstructed image is expected to be similar to the original test image. On the other hand, if the image contains anomalous regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. Therefore, the anomalous regions can be detected by comparing the original test image and its reconstructed one.
Problem formulation.

### v4

# Quantifying Statistical Significance in Diffusion-Based Anomaly Localization via Selective Inference

@@ Abstract
Anomaly localization in images—identifying regions that deviate from expected patterns—is vital in applications such as medical diagnosis and industrial inspection. A recent trend is the use of image generation models in anomaly localization, where these models generate normal-looking counterparts of anomalous images, thereby allowing flexible and adaptive anomaly localization. However, these methods inherit the uncertainty and bias implicitly embedded in the employed generative model, raising concerns about the reliability. To address this, we propose a statistical framework based on selective inference to quantify the significance of detected anomalous regions. Our method provides $p$-values to assess the false positive detection rates, providing a principled measure of reliability. As a proof of concept, we consider anomaly localization using a diffusion model and its applications to medical diagnoses and industrial inspections. The results indicate that the proposed method effectively controls the risk of false positive detection, supporting its use in high-stakes decision-making tasks.
@@ 1 Introduction
Anomaly localization using image generation models, particularly diffusion models, has shown great promise across diverse domains such as medical diagnosis and industrial inspection Li et al. (2023); Iqbal et al. (2023); Lu et al. (2023); Zhang et al. (2023); Fontanella et al. (2024); Tebbe and Tayyub (2024); Sheng et al. (2024). These models reconstruct a normal-looking version of an input image, and differences between the input and the reconstruction highlight potential anomalies. Compared to traditional methods, generative approaches are highly suitable for settings where annotations for anomalous regions are unavailable. Moreover, generative approaches can flexibly handle heterogeneity by adapting to individual images–e.g., patient-specific characteristics in medical diagnosis and product-specific traits in industrial inspection. Among various image generation models, diffusion models, in particular, offer high fidelity and stability, outperforming other methods in image quality and anomaly localization.
While generative approaches offer powerful and flexible capabilities for anomaly localization, a major concern is that the inherent uncertainty and bias in generative models can affect localization performance Fithian et al. (2014); Taylor and Tibshirani (2015); Lee et al. (2016); Duy and Takeuchi (2022); Miwa et al. (2023); Shiraishi et al. (2024). These models are trained on specific datasets composed of normal images, and the quality of the generated normal-looking images depends heavily on the distribution of the dataset and how well the model has learned the underlying distribution. As a result, uncertainties or biases in the dataset or training process can cause incorrect reconstructions, leading to inaccurate localizations and misidentification of anomalies. Such risks are especially critical in high-stakes domains such as medical diagnosis and industrial inspection, where even minor errors can have serious consequences. Therefore, it is essential to incorporate rigorous uncertainty quantification framework and statistical safeguards to ensure reliable deployment in critical applications.
To address this issue, we propose a statistical testing framework based on Selective Inference (SI) to assess the statistical significance of detected anomalies. SI has recently emerged as a promising approach for conducting statistical inference on hypotheses that are selected based on observed data. In this framework, inference is performed using the conditional distribution given the selection event, thereby accounting for the uncertainty and bias associated with the hypothesis selection. Following this SI framework, our key idea is to apply statistical testing to detected anomalies, conditioned on the fact that the anomalous regions were identified using a specific generative model. This allows us to quantify the statistical significance of detected anomalies as a valid $p$-value, providing a rigorous estimate of the false positive rate and offering a principled metric for reliability.
In this paper, as a proof of concept for the proposed statistical testing framework, we focus on a standard denoising diffusion probabilistic model (DDPM) Ho et al. (2020); Song et al. (2022) among various diffusion-based anomaly detection (AD) methods , and its applications to medical diagnostics and industrial inspections. However, the proposed framework is readily generalizable to a broader range of diffusion model architectures and is applicable to semi-supervised AD problems in other domains.
Related works
@@
Diffusion models have been effectively utilized in anomaly localization problems Pinaya et al. (2022); Fontanella et al. (2024); Wyatt et al. (2022); Mousakhan et al. (2023). In this context, the DDPM is commonly used Ho et al. (2020); Song et al. (2022). During the training phase, a DDPM model learns the distribution of normal medical images by iteratively adding and then removing noise. In the test phase, the model attempts to reconstruct a new test image. If the image contains anomalous regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. The discrepancies between the original and the reconstructed image are then analyzed to identify and highlight anomalous regions. Other types of generative AI has also been used for anomalous region detection task Baur et al. (2021); Chen and Konukoglu (2018); Chow et al. (2020); Jana et al. (2022).
@@
SI was first introduced within the context of reliability evaluation for linear model features when they were selected using a feature selection algorithm Lee and Taylor (2014); Lee et al. (2016); Tibshirani et al. (2016), and then extended to more complex feature selection methods Yang et al. (2016); Suzumura et al. (2017); Hyun et al. (2018); Rügamer and Greven (2020); Das et al. (2021). Then, SI proves valuable not only for feature selection problems but also for statistical inference across various data-driven hypotheses, including unsupervised learning tasks Chen and Bien (2020); Tsukurimichi et al. (2021); Tanizaki et al. (2020); Duy et al. (2022); Le Duy et al. (2024); Lee et al. (2015); Gao et al. (2022); Duy et al. (2020); Jewell et al. (2022). The fundamental idea of SI is to perform an inference conditional on the hypothesis selection event, which mitigates the selection bias issue even when the hypothesis is selected and tested using the same data. To conduct SI, it is necessary to derive the sampling distribution of test statistic conditional on the hypothesis selection event. To the best of our knowledge, SI was applied to statistical inferences on several deep learning models Duy et al. (2022); Miwa et al. (2023); Shiraishi et al. (2024); Miwa et al. (2024), but none of them works on image generation by diffusion models.
Our contributions
The main contributions of our study are summarized as follows.  [^1] We note that our contribution is not the development of a new diffusion-based anomaly localization algorithm, but rather the introduction of a rigorous statistical testing framework designed to quantify the statistical reliability of anomalous regions identified by diffusion-based AD methods.
  • We propose a novel statistical testing framework to assess the significance of anomaly localization results derived from diffusion model-based methods, offering a principled basis for evaluating the reliability of detected anomalies.
  • We implement the SI framework for diffusion models by deriving the sampling distribution conditional on the selection event induced by the diffusion model, which requires developing non-trivial computational techniques tailored to the generative sampling process.
  • We provide theoretical justification for the proposed method and validate its effectiveness through extensive numerical experiments in medical diagnosis and industrial inspection scenarios. The results highlight the robustness and practical utility of our method.
The implementation code for reproducing all experimental results is provided at https://github.com/tkatsuoka/DAL-test
Figure 1: Schematic illustration of anomaly localization on a brain MRI image dataset using a diffusion model and the proposed DAL-Test. When a test image—potentially containing an anomalous region—is fed into a trained diffusion model, its normal-looking version is generated through the forward and reverse processes. By comparing the input image with the generated normal-looking version, the anomalous region can be identified. We propose the Diffusion-based Anomaly Localization (DAL) Test, which leverages the selective inference framework to compute valid $p$-values that quantify the statistical significance of anomalous regions detected by a diffusion model, based on a test statistic defined over the input and reference images.
2 Anomaly localization by diffusion models
This section describes the anomaly localization task based on a diffusion model, which is explored as a proof of concept in this study. The process of anomaly localization using generative models can generally be divided into two phases. First, during the training phase, a denoising diffusion probabilistic model (DDPM) is trained using a dataset composed exclusively of normal images. The model learns the distribution of normal images through two key processes: the diffusion process, in which noise is gradually added to an image, and the reverse diffusion process, in which the original image is reconstructed from noise. Through this procedure, the model enhances its capacity to reconstruct normal image structures by acquiring denoising capabilities at each step. Next, during the testing phase, the reverse diffusion process is conditionally applied to an unseen input test image. In this step, the model reconstructs an image that closely resembles the input but conforms to its learned notion of “normality”, causing anomalous regions to be poorly reproduced. An anomaly score is then computed based on the difference between the reconstructed image and the input test image, and the spatial distribution of this score is analyzed to localize anomalies. By applying thresholding to the score map, anomalous regions can be clearly identified.
@@ 2 Anomaly localization by diffusion models
In this study, for the purpose of proof of concept, we adopt standard denoising diffusion models as our choice of diffusion model Ho et al. (2020); Song et al. (2022). The following outlines the image reconstruction process of a trained DDPM. Given a test image which possibly contain anomalous regions, a denoising diffusion model is used to generate the corresponding normal image. The reconstruction process consists of two processes called forward process (or diffusion process) and reverse process. In the forward process, noise is sequentially added to the test image so that it converges to a standard Gaussian distribution $\mathcal{N}(\bm{0},I)$. Let $\bm{x}$ be an image represented as a vector with each element corresponding to a pixel value. Given an original test image $\bm{x}_{0}$, noisy images $\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{T}$ are sequentially generated, where $T$ is the number of noise addition steps. We consider the distribution of the original and noisy test images, which is denoted by $q(\bm{x})$, and approximate the distribution by a parametric model $p_{\theta}(\bm{x})$ with $\theta$ being the parameters. Using a sequence of noise scheduling parameters $0<\beta_{1}<\beta_{2},<\dots<\beta_{T}<1$, the forward process is written as
$q(\bm{x}_{1:T}|\bm{x}_{0}):=\prod_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1}),\;\;\text{where}\;\;q(\bm{x}_{t}|\bm{x}_{t-1}):=\mathcal{N}(\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}I).$ (1)
By the reproducibility of the Gaussian distribution, $\bm{x}_{t}$ can be rewritten by a linear combination of $\bm{x}_{0}$ and $\epsilon_{t}$, i.e.,
$\bm{x}_{t}=\sqrt{\alpha_{t}}\bm{x}_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t},\;\;\text{with}\;\;\epsilon_{t}\sim\mathcal{N}(\mathbf{0},I),$ (2)
@@
where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$. In the reverse process, a parametric model in the form of $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1};\mu_{\theta}(\bm{x}_{t},t),\beta_{t}I)$ is employed, where $\mu_{\theta}(\bm{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\bm{x}_{t})$. Typically, a U-Net is used as the model architecture for $\epsilon^{(t)}_{\theta}(\bm{x}_{t})$. In DDPM Ho et al. (2020), the loss function for training the noise component is simply written as $||\epsilon^{(t)}_{\theta}(\bm{x}_{t})-\epsilon_{t}||^{2}_{2}$. Based on (2), given a noisy image $\bm{x}_{t}$ after $t$ steps, the reconstruction of the image in the previous step $\bm{x}_{t-1}$ is obtained as
$\bm{x}_{t-1}=\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\bm{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{(t)}(\bm{x}_{t})+\sigma_{t}\epsilon_{t},$ (3)
@@
$f^{(t)}_{\theta}(\bm{x}_{t}):=(\bm{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t}))/\sqrt{\alpha_{t}},$ (4)
@@
Here, $\eta$ is a hyperparameter that controls the randomness in the reverse process. By setting $\eta=1$, we can create new images by stochastic sampling. On the other hand, if we set $\eta=0$, deterministic sampling is used for image generation. By recursively sampling as in (3), we can obtain a reconstructed image of the original input $\bm{x}_{0}$.
In practice, the reverse process starts from $\bm{x}_{T^{\prime}}$ with $T^{\prime}<T$. Namely, we reconstruct the original input image not from the completely noisy one, but from a one which still contains individual information of the original input image. The smaller $T^{\prime}$ ensures that the reconstructed image preserves fine details of the input image. Conversely, the larger $T^{\prime}$ results in the retention of only large scale features, thereby converting more of the anomalous regions into normal regions Ho et al. (2020); Mousakhan et al. (2023).
@@
0:  Input image $\bm{x}$
1:  $\bm{x}_{T^{\prime}}\leftarrow\sqrt{\alpha_{T^{\prime}}}\bm{x}+\sqrt{1-\alpha_{T^{\prime}}}\epsilon$
@@
3:     $f^{(t)}_{\theta}(\bm{x}_{t})\leftarrow(\bm{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t}))/\sqrt{\alpha_{t}}$
4:     $\bm{x}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\bm{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t})+\sigma_{t}\epsilon_{t}$
@@
5:  Reconstructed image $\bm{x}_{0}$
3 Statistical test for diffusion-based anomalous localization
In this section, we formulate the statistical test for detecting anomalous regions by a trained diffusion model. As shown in Figure 1, anomalous region detection by diffusion models is conducted as follows. First, in the training phase, the diffusion model is trained only on normal images. Then, in the test phase, we feed a test image which might contain anomalous regions into the trained diffusion model, and reconstruct it back from a noisy image $\mathbf{x}_{T^{\prime}}$ at step $T^{\prime}<T$. By appropriately selecting $T^{\prime}$, we can generate a normal image that retain individual characteristics of the test input image. If the image does not contain anomalous regions, the reconstructed image is expected to be similar to the original test image. On the other hand, if the image contains anomalous regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. Therefore, the anomalous regions can be detected by comparing the original test image and its reconstructed one.
Problem formulation